OpenStudy (themathasker):

Jerome has 1/2 of the groups video games at his house. Mario have 2/5 of the video games at his house. What fraction of the group's video games are either at Jerome or Mario's house?

OpenStudy (ajspeller):

1/2 = 50% 2/5 = 40% 50% + 40% = 90% 90% = 9/10 = 0.9